OpenStudy (anonymous):

In the TV show "Breaking Bad" the characters attempt to use HF acid to dissolve guns (among other things). Here we consider instead dissolving guns (which we will assume are pure iron) with sulfuric acid. Complete the balanced reaction for reacting iron in dilute sulfuric acid to form aqueous FeSO4. Do not worry about formatting subscripts (i.e. O2 to represent diatomic oxygen gas is fine). Fe + H2SO4 → FeSO4 + __H2(g)___

OpenStudy (anonymous):

How many liters of 1 molar sulfuric acid would be required to dissolve 1 kg of iron? Assume the reaction from the previous part goes to completion. The molecular mass of Fe is 55.85 g/mol. ANSWER: ??????????????????????? It is apparent that they must use an outside source of electrical power to drive the dissolution of the iron. They use a small current so that the dissolution proceeds with the minimum voltage required. Assume standard values for the electrochemical potentials. How much electrical energy supplied this way is thus required to dissolve an additional 1 kg of iron? Give your answer in kJ. ANSWER: ???????????????????????????? The characters realize that their hideout has been discovered by the police and they still have a last handgun that weighs 0.25 kg to dissolve. The cops will get there in an hour, so they have to speed up the reaction, by driving it at a higher current. What is the minimum total voltage in volts they'll need to drive the reaction at to get rid of the gun in time? Consider excess potential because of activation losses only and the exchange current I0 to be 1 A for the reaction over the surface of the entire tank (not a current density). α is 0.5 and everything is being done at room temperature. Assume standard electrochemical potentials. ANSWER: ???????????????????????????

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

1st question: 17.9

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

@dan815

OpenStudy (aaronq):

For the second, use: \(Fe^{2+} + 2 e^- \rightarrow Fe\) \(E_{red}\)= -0.44 V standard potentials are at 1M of 1 L, so 1 mol, so: so find moles of Fe (equivalent to 1kg), then multiply by the oxidation potential (+0.44V). For the 3rd: i'm not sure what \(\alpha\) is, do you know what equation you're supposed to use?

OpenStudy (anonymous):

@aaronq can u solve it

OpenStudy (aaronq):

nope. you should do this yourself, we're here to help not do it for you.

OpenStudy (anonymous):

i got the calculations but i dont got the right one

OpenStudy (aaronq):

post what you've done then, we'll check it

OpenStudy (anonymous):

@aaronq dont think calculate dcorrectly....

OpenStudy (anonymous):

.44 x 1000 x 33

OpenStudy (anonymous):

i dont think that is right

OpenStudy (aaronq):

units?

OpenStudy (anonymous):

if i look at yours it doesn't work but i used two of what u listed but the other one is g

OpenStudy (anonymous):

Fe molecular weight is 55.85

OpenStudy (anonymous):

can you lay out wher to plug in? the equation?

OpenStudy (aaronq):

it should be \(0.44 ~V/mol*n_{Fe}\) so \(\dfrac{0.44 V/mol*1000g}{55.85 g/mol)}\)

OpenStudy (anonymous):

7.878245299910474485228290062667860340197

OpenStudy (anonymous):

is that what u listed?

OpenStudy (anonymous):

i dont thin it is correct

OpenStudy (anonymous):

@aaronq doesn't look right

OpenStudy (aaronq):

then you need to find the energy using: V = E/ Q Q = electron = 1.60217657 × 10^-19 coulombs

OpenStudy (anonymous):

so use .44?

OpenStudy (anonymous):

electron x .44?

OpenStudy (aaronq):

use the value you got after

OpenStudy (aaronq):

7.878

OpenStudy (anonymous):

so the long answer divided by eletron?

OpenStudy (aaronq):

no, you're finding E, so E=V*Q

OpenStudy (anonymous):

if u multiply em you get: 0.000000000000000001262234003222918531781557743957027752909586258429

OpenStudy (anonymous):

is that what you got?

OpenStudy (anonymous):

@aaronq did we get the same answer?

OpenStudy (aaronq):

hm its a very small number.. but it looks right.

OpenStudy (anonymous):

let me double check

OpenStudy (anonymous):

is that in KJ?

OpenStudy (aaronq):

no that would be in J

OpenStudy (anonymous):

conver tto KJ

OpenStudy (anonymous):

times 1000?

OpenStudy (aaronq):

divide

OpenStudy (anonymous):

if you divide it it become: 0.00000000000000000000126223400322291853178155774395702775291

OpenStudy (anonymous):

i think it became smaller....

OpenStudy (aaronq):

yeah it gets smaller when you go from a smaller unit to a larger unit

OpenStudy (anonymous):

but both answer are wrong....

OpenStudy (anonymous):

did you try to calculate it yourself?

OpenStudy (anonymous):

to get KJ in the end

OpenStudy (aaronq):

the first one is simple stoichiometry: Fe + H2SO4 → FeSO4 + __H2(g)__ 1 kg of Fe to moles moles of Fe = moles of H2SO4 L of solution= moles of H2SO4/Molarity

OpenStudy (anonymous):

i didnt convert Fe to KG.....

OpenStudy (anonymous):

omg.....

OpenStudy (aaronq):

you're supposed to use it in grams, because the molar mass is in grams per mole

OpenStudy (anonymous):

can you map out the equation with the numbers in a better way?

OpenStudy (anonymous):

even your setup we get wrong answer

OpenStudy (anonymous):

im gona rage quit lol

OpenStudy (aaronq):

moles of iron= 1000g/55.845g/mol \(L_{H_2SO_4}=\dfrac{\dfrac{1000g}{55.845g/mol}}{1~ mol/L}\)

OpenStudy (aaronq):

so the answer the previous person gave was right.

OpenStudy (anonymous):

wrong.....it says

OpenStudy (anonymous):

maybe gotta convert to one decimal format

OpenStudy (anonymous):

i dont know anymore

OpenStudy (aaronq):

try 18 L

OpenStudy (anonymous):

for the answer or for the number to calucaulte?

OpenStudy (aaronq):

the answer, just rounded up

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

i dont kthink u should round it...

OpenStudy (anonymous):

needs a raw answer in kJ

OpenStudy (aaronq):

for the first question? it's asking for liters

OpenStudy (anonymous):

the second question

OpenStudy (anonymous):

someone already gave the first answer " 1st question: 17.9 "

OpenStudy (anonymous):

if you read above your post someone answered it already we need the second and third answers

OpenStudy (anonymous):

@aaronq we need second and third but you said u dont know third one so we do the second one

OpenStudy (aaronq):

i know that. you said it was wrong, so i was checking it

OpenStudy (anonymous):

jus raw answer is good enough no round

OpenStudy (aaronq):

lol thats not what you said, but whatever.

OpenStudy (anonymous):

we just need second question answer

OpenStudy (aaronq):

hm it says "an additional kg of iron".. so maybe multiply that number by 2, because we only did it for 1 kg

OpenStudy (anonymous):

ok what did u get then?

OpenStudy (aaronq):

im not doing the calculations

OpenStudy (anonymous):

im already lost...i lost all my numbers

OpenStudy (anonymous):

i rage quitted my calculator

OpenStudy (aaronq):

go back up, they're somewhere on the page

OpenStudy (anonymous):

ok so give me the numebr to calucator x 2

OpenStudy (anonymous):

which number do i multiply by 2

OpenStudy (aaronq):

the one we found earlier

OpenStudy (anonymous):

ill caulaate it jus give me the number

OpenStudy (aaronq):

its up in this page, look for it

OpenStudy (anonymous):

0.00000000000000000000252446800644583706356311548791405550582 answer is this? can u check ur answer too?i already multipolied it

OpenStudy (anonymous):

OMG WRONG!!!!

OpenStudy (anonymous):

I RAGE QUIT OMGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGG

OpenStudy (aaronq):

relax man

OpenStudy (anonymous):

GONA CUT OF MY HAIR NOW ALL OF IT!

OpenStudy (aaronq):

lol

OpenStudy (anonymous):

i like few numbers but u give me too many numbers and that has so many zeros i cant handle it

OpenStudy (anonymous):

what answer did you get jus post it

OpenStudy (anonymous):

need to system check it if it works

OpenStudy (aaronq):

2.5*10^-21

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

its over....game over man

OpenStudy (anonymous):

all the answer give X which is wrong